OpenStudy (anonymous):

find each function value, if it exists. f(t)=√(t^2+1) A.) f(-8)= ? (Simplify your answer. Type an exact answer, using radicals as needed.) B.) The square root is not a real number.

OpenStudy (anonymous):

you already asked this quetion? or it was somene else like half hour ago?

OpenStudy (anonymous):

yeah it was you.. so why you asking it again?

OpenStudy (anonymous):

a certain sense of deja vu

OpenStudy (anonymous):

\[f(-8)=\sqrt{(-8)^2+1}=\sqrt{65}\]

OpenStudy (anonymous):

cuzz know one would give me an answer!!!

OpenStudy (anonymous):

SO... f(-8)= ?